OpenStudy (anonymous):

A car accelerates from rest at -3.00 m/s^s

OpenStudy (anonymous):

what are you finding?

OpenStudy (anonymous):

I'm trying to find the velocity after 5.0 seconds and the displacement after the same time

OpenStudy (anonymous):

which physics are you taking?

OpenStudy (anonymous):

Honors Physics

OpenStudy (anonymous):

do you use calculus in it?

OpenStudy (anonymous):

if not, v = vi + at

OpenStudy (anonymous):

vi is initial velocity, which is zero because it starts from rest

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

x = xi + vi t + 1/2 at^2

OpenStudy (anonymous):

where xi is initial position and is zero

OpenStudy (anonymous):

vi is also zero

OpenStudy (anonymous):

So 0+0+1/2 at^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just gave you those equations because they are standard physics equations

OpenStudy (anonymous):

and are applicable to almost every kinematics situation

OpenStudy (anonymous):

sometimes the objects don't start at rest or start in another place another good one to know is vf^2 = vi^2 + 2a(xf - xi) vf is final velocity, vi is initial velocity, xf is final position, and xi is initial position

OpenStudy (anonymous):

at would be the 5 seconds?

OpenStudy (anonymous):

t = 5 so (.5) * (-3) * (5^2) is that what you meant?

OpenStudy (anonymous):

Yes